OpenStudy (anonymous):

What are the solutions to the equation x² = 625? Choose all answers that are correct. A. –312.5 B. –25 C. 25 D. 312.5 I'm pretty sure i know what the answer's are but i just want to make sure

OpenStudy (anonymous):

How do you think you would go about solving this?

OpenStudy (anonymous):

i just do 625/25 or 625/-25 and the answer's are 25 and -25 so that's how i get my answer's but i may be wrong

OpenStudy (jdoe0001):

\(\bf x^{\color{brown}{ 2}}=625\implies \sqrt[{\color{brown}{ 2}}]{x^{\color{brown}{ 2}}}=\sqrt[{\color{brown}{ 2}}]{625}\implies x=\sqrt[{\color{brown}{ 2}}]{625}\)

OpenStudy (anonymous):

That would be a way of checking if a guess was right, but that only works if they give you options to choose from. To solve, you would take the square root of both sides. Some numbers we just know are perfect squares, like 4, 9, 16 etc. But for larger numbers we would use a calculator if possible. We know that x will have two answers, a positive and negative answer. This is because both the positive and negative answers squared (multiplied by themselves ) will give a positive answer. So take the square root of 625 and it would be both plus or minus this number.